# Continuous Invariance Learning

### v1

@@ Abstract
Invariance learning methods aim to learn invariant features in the hope that they generalize under distributional shifts. Although many tasks are naturally characterized by continuous domains, current invariance learning techniques generally assume categorically indexed domains. For example, auto-scaling in cloud computing often needs a CPU utilization prediction model that generalizes across different times (e.g., time of a day and date of a year), where ‘time’ is a continuous domain index. In this paper, we start by theoretically showing that existing invariance learning methods can fail for continuous domain problems. Specifically, the naive solution of splitting continuous domains into discrete ones ignores the underlying relationship among domains, and therefore potentially leads to suboptimal performance. To address this challenge, we then propose Continuous Invariance Learning (CIL), which extracts invariant features across continuously indexed domains. CIL is a novel adversarial procedure that measures and controls the conditional independence between the labels and continuous domain indices given the extracted features. Our theoretical analysis demonstrates the superiority of CIL over existing invariance learning methods. Empirical results on both synthetic and real-world datasets (including data collected from production systems) show that CIL consistently outperforms strong baselines among all the tasks.
@@
Keywords Invariance Learning  $\cdot$ Domain generalization  $\cdot$ Continuous Domains
@@ 1 Introduction
Machine learning models have shown impressive success in many applications, including computer vision He et al. (2016), natural language processing Liu et al. (2019), speech recognition Deng et al. (2013), etc. These models normally take the independent identically distributed (IID) assumption and assume that the testing samples are drawn from the same distribution as the training samples. However, these assumptions can be easily violated when there are distributional shifts in the testing datasets. This is also known as the out-of-distribution (OOD) generalization problem.
@@
Consider invariant features that can generalize under distributional shift and spurious features that are unstable. Let $\boldsymbol{x}$ denote the input, $\boldsymbol{y}$ denote the output, $\boldsymbol{t}$ denote the domain index, and $\mathbb{E}^{t}$ denote taking expectation in domain $\boldsymbol{t}$. Consider the model composed of a featurizer $\Phi$ and a classifier $\boldsymbol{w}$ (Arjovsky et al., 2019). Invariant Risk Minimization (IRM) proposes to learn $\Phi(\boldsymbol{x})$ by merely using invariant features, which yields the same conditional distribution $\mathbb{P}^{t}(\boldsymbol{y}|\Phi(\boldsymbol{x}))$ across all domains $\boldsymbol{t}$, i.e., $\mathbb{P}(\boldsymbol{y}|\Phi(\boldsymbol{x}))=\mathbb{P}(\boldsymbol{y}|\Phi(\boldsymbol{x}),\boldsymbol{t})$, which is equivalent to $\boldsymbol{y}\perp\boldsymbol{t}|\Phi(\boldsymbol{x})$ (Arjovsky et al., 2019). Existing approximation methods propose to learn a featurizer by matching $\mathbb{E}^{t}[\boldsymbol{y}|\Phi(\boldsymbol{x})]$ for different $\boldsymbol{t}$ (See Section 2.1) for details.).
@@
A case study on REx. However, in the continuous environment setting, there are only very limited samples in each environment. So the finite sample estimations $\hat{\mathbb{E}}^{t}[\boldsymbol{y}|\Phi(\boldsymbol{x})]$ can deviate significantly from the expectation $\mathbb{E}^{t}[\boldsymbol{y}|\Phi(\boldsymbol{x})]$. In Section 2.2, We conduct a theoretical analysis of REx (Krueger et al., 2021), a popular variant of IRM. Our analysis shows that when there is a large number of domains and limited sample sizes per domain (i.e., in the continuous domain tasks), REx fails to identify invariant features with a constant probability. This is in contrast to the results in discrete domain tasks, where REx can reliably identify invariant features given a sufficient number of samples in each discrete domain.
@@
The generality of our results. Other popular approximations of IRM also suffer from this limitation, such as IRMv1 (Arjovsky et al., 2019), REx (Krueger et al., 2021), IB-IRM (Ahuja et al., 2021), IRMx Chen et al. (2022), IGA Koyama and Yamaguchi (2020), BIRM Lin et al. (2022a), IRM Game Ahuja et al. (2020), Fisher Rame et al. (2022) and SparseIRM (Zhou et al., 2022). They employ positive losses on each environment $\boldsymbol{t}$ to assess whether $\hat{\mathbb{E}}^{t}[\boldsymbol{y}|\Phi(\boldsymbol{x})]$ matches the other environments. However, since the estimations $\hat{\mathbb{E}}^{t}[\boldsymbol{y}|\Phi(\boldsymbol{x})]$ are inaccurate, these methods fail to identify invariant features. We then empirically verify the ineffectiveness of these popular invariance learning methods in handling continuous domain problems. One potential naive solution is to split continuous domains into discrete ones. However, this splitting scheme ignores the underlying relationship among domains and therefore can lead to sub-optimal performance, which is further empirically verified in Section 2.2
@@
Figure 1: Illustration of distributional shifts in discrete and continuous domains. Existing IRM methods focus on problems with discrete domains, which is inconsistent with many real-world tasks. Our work therefore aims to extend IRM to continuous domains.
@@
Our methods. Recall that our task is to learn a feature extractor $\Phi(\cdot)$ that extracts invariant features from $\boldsymbol{x}$ into invariant features, i.e., learn $\Phi$ which elicits $\boldsymbol{y}\perp\boldsymbol{t}|\Phi(\boldsymbol{x})$. Previous analysis shows that it is infeasible to align $\mathbb{E}^{t}[\boldsymbol{y}|\Phi(\boldsymbol{x})]$ among different domain $\boldsymbol{t}$ in continuous domain tasks. Instead, we propose to align $\mathbb{E}^{y}[\boldsymbol{t}|\Phi(\boldsymbol{x})]$ for each class $\boldsymbol{y}$. We can start by training two domain regressors, i.e., $h(\Phi(\boldsymbol{x}))$ and $g(\Phi(\boldsymbol{x}),\boldsymbol{y})$, to regress over the continuous domain index $\boldsymbol{t}$ using L1 or L2 loss. Notably, the L1/L2 distance between the predicted and ground-truth domains captures the underlying relationship between continuous domain indices and can lead to similar domain index prediction losses between domains on both $h$ and $g$ if $\Phi(\boldsymbol{x})$ only extracts invariant features. Specifically, the loss achieved by $h$, i.e., $\mathbb{E}\|\boldsymbol{t}-h(\Phi(\boldsymbol{x}))\|_{2}^{2}$, and the loss achieved by $g$, i.e., $\mathbb{E}\|\boldsymbol{t}-g(\Phi(\boldsymbol{x}),\boldsymbol{y})\|_{2}^{2}$ would be the same if $\Phi(\boldsymbol{x})$ only extracts invariant features. The whole procedure is then formulated into a mini-max framework as shown in Section 3.
@@
In Section 3.1, we emphasize the theoretical advantages of CIL over existing IRM approximation methods in continuous domain tasks. The limited number of samples in each domain makes it challenging to obtain an accurate estimation for $\mathbb{E}^{t}[\boldsymbol{y}|\Phi(\boldsymbol{x})]$, leading to the ineffectiveness of existing methods. However, CIL does not suffer from this problem because it aims to align $\mathbb{E}^{y}[\boldsymbol{t}|\Phi(\boldsymbol{x})]$, which can be accurately estimated due to the ample number of samples available in each class (also see Appendix D for the discussion of the relationship between $\mathbb{E}^{t}[\boldsymbol{y}|\Phi(\boldsymbol{x})]$ and $\mathbb{E}^{y}[\boldsymbol{t}|\Phi(\boldsymbol{x})]$). In Section 4, we carry out experiments on both synthetic and real-world datasets (including data collected from production systems in Alipay and vision datasets from Wilds-time (Yao et al., 2022)). CIL consistently improves over existing invariance learning baselines in all the tasks. We summarize our contributions as follows:
@@
  • We identify the problem of learning invariant features under continuous domains and theoretically demonstrate that existing methods which treat domain as categorical index can fail with constant probability, which is further verified by empirical study.
@@
  • We propose Continuous Invariance Learning (CIL) as the first general training framework to address this problem. We also show theoretical advantages of CIL over existing IRM approximation methods on continuous domain tasks.
@@ 2.1 Preliminaries
Notations. Consider the input and output $(\boldsymbol{x},\boldsymbol{y})$ in the space $\mathcal{X}\times\mathcal{Y}$. Our task is to learn a function $f_{\theta}\in\mathcal{F}:\mathcal{X}\xrightarrow[]{}\mathcal{Y}$, parameterized by $\theta$. We further denote the domain index as $\boldsymbol{t}\in\mathcal{T}$. We have $(\boldsymbol{x},\boldsymbol{y},\boldsymbol{t})\sim\mathbb{P}(\boldsymbol{x},\boldsymbol{y}|\boldsymbol{t})\times\mathbb{P}(\boldsymbol{t})$. Denote the training dataset as $\mathcal{D}^{\text{tr}}:=\{(\mathbf{x}_{i}^{\text{tr}},\mathbf{y}_{i}^{\text{tr}},\mathbf{t}_{i}^{\text{tr}})\}_{i=1}^{n}$. The goal of domain generalization is to learn a function $f_{\theta}$ on $\mathcal{D}^{\text{tr}}$ that can perform well in unseen testing dataset $\mathcal{D}^{\text{te}}$. Since the testing domain $\boldsymbol{t}^{\text{te}}$ differs from $\boldsymbol{t}^{\text{tr}}$, $\mathbb{P}(\boldsymbol{x},\boldsymbol{y}|\boldsymbol{t}^{\text{te}})$ is also different from $\mathbb{P}(\boldsymbol{x},\boldsymbol{y}|\boldsymbol{t}^{\text{tr}})$ due to distributional shift. Following (Arjovsky et al., 2019), we assume that $\boldsymbol{x}$ is generated from invariant feature $\boldsymbol{x}_{v}$ and spurious feature $\boldsymbol{x}_{s}$ by some unknown function $\xi$, i.e., $\boldsymbol{x}=\xi(\boldsymbol{x}_{v},\boldsymbol{x}_{s})$. By the invariance property (more details shown in Appendix C), we have $\boldsymbol{y}\perp\boldsymbol{t}|\boldsymbol{x}_{v}$ for all $\boldsymbol{t}\in\mathcal{T}$. The target of invariance learning is to make $f_{\theta}$ only dependent on $\boldsymbol{x}_{v}$.
@@
Invariance Learning. To learn invariant features, Invariant Risk Minimization (IRM) first divides the neural network $\theta$ into two parts, i.e., the feature extractor $\Phi$ and the classifier $\boldsymbol{w}$ (Arjovsky et al., 2019). The goal of invariance learning is to extracted $\Phi(\boldsymbol{x})$ which satisfies $\boldsymbol{y}\perp\boldsymbol{t}|\Phi(\boldsymbol{x})$.
@@
Existing Approximation Methods. Suppose we have collected the data from a set of discrete domains, $\mathcal{T}=\{\boldsymbol{t}_{1},\boldsymbol{t}_{2},...,\boldsymbol{t}_{M}\}$. The loss in domain $\boldsymbol{t}$ is $\mathcal{R}^{t}(\boldsymbol{w},\Phi):=\mathbb{E}^{t}[\ell(\boldsymbol{w}(\Phi(\boldsymbol{x})),\boldsymbol{y})]$ Since is hard to validate $\boldsymbol{y}\perp\boldsymbol{t}|\Phi(\boldsymbol{x})$ in practice, existing works propose to align $\mathbb{E}^{t}[\boldsymbol{y}|\Phi(\boldsymbol{x})]$ for all $t$ as an approximation. Specifically, if $\Phi(\boldsymbol{x})$ merely extract invariant features, $\mathbb{E}^{t}[\boldsymbol{y}|\Phi(\boldsymbol{x})]$ would be the same in all $\boldsymbol{t}$. Let $\boldsymbol{w}^{t}$ denote the optimal classifier for domain $\boldsymbol{t}$, i.e., $\boldsymbol{w}^{t}:=\operatorname*{arg\,min}_{\boldsymbol{w}}\mathcal{R}^{t}(\boldsymbol{w},\Phi)$. We have $\boldsymbol{w}^{t}(\Phi(\boldsymbol{x}))=\mathbb{E}^{t}[\boldsymbol{y}|\Phi(\boldsymbol{x})]$ hold if the function space of $\boldsymbol{w}$ is sufficiently large Li et al. (2022). So if $\Phi(\boldsymbol{x})$ relies on invariant features, $\boldsymbol{w}^{t}$ would be the same in all $\boldsymbol{t}$. Existing approximation methods try to ensure $\boldsymbol{w}^{t}$ that is the same for all environments to align $\mathbb{E}^{t}[\boldsymbol{y}|\Phi(\boldsymbol{x})]$ (Arjovsky et al., 2019; Lin et al., 2022a; Ahuja et al., 2021). Further variants also include checking whether $\mathcal{R}^{t}(\boldsymbol{w},\Phi)$ is the same (Krueger et al., 2021; Ahuja et al., 2020; Zhou et al., 2022; Chen et al., 2022), or the gradient is the same for all $\boldsymbol{t}$ (Rame et al., 2022; Koyama and Yamaguchi, 2020). For example, REx (Krueger et al., 2021) penalizes the variance of the loss in domains:
$\mathcal{L}_{\text{REx}}=\sum\nolimits_{t}\mathcal{R}^{t}(\boldsymbol{w},\Phi)+\lambda\text{Var}(\mathcal{R}^{t}(\boldsymbol{w},\Phi)).$ (1)
where $\text{Var}(\mathcal{R}^{t}(\boldsymbol{w},\Phi))$ is the variance of the losses in among domains.

### v2

@@ Abstract
Invariance learning methods aim to learn invariant features in the hope that they generalize under distributional shifts. Although many tasks are naturally characterized by continuous domains, current invariance learning techniques generally assume categorically indexed domains. For example, auto-scaling in cloud computing often needs a CPU utilization prediction model that generalizes across different times (e.g., time of a day and date of a year), where ‘time’ is a continuous domain index. In this paper, we start by theoretically showing that existing invariance learning methods can fail for continuous domain problems. Specifically, the naive solution of splitting continuous domains into discrete ones ignores the underlying relationship among domains, and therefore potentially leads to suboptimal performance. To address this challenge, we propose Continuous Invariance Learning (CIL), which extracts invariant features across continuously indexed domains. CIL is a novel adversarial procedure that measures and controls the conditional independence between the labels and continuous domain indices given the extracted features. Our theoretical analysis demonstrates the superiority of CIL over existing invariance learning methods. Empirical results on both synthetic and real-world datasets (including data collected from production systems) show that CIL consistently outperforms strong baselines among all the tasks.
@@ 1 Introduction
Machine learning models have shown impressive success in many applications, including computer vision (He et al., 2016), natural language processing (Liu et al., 2019), speech recognition (Deng et al., 2013), etc. These models normally take the independent identically distributed (IID) assumption and assume that the testing samples are drawn from the same distribution as the training samples. However, these assumptions can be easily violated when there are distributional shifts in the testing datasets. This is also known as the out-of-distribution (OOD) generalization problem.
@@
Consider invariant features that can be generalized under distributional shift and spurious features that are unstable. Let $\bm{x}$ denote the input, $\bm{y}$ denote the output, $\bm{t}$ denote the domain index, and $\mathbb{E}^{t}$ denote taking expectation in domain $\bm{t}$. Consider the model composed of a featurizer $\Phi$ and a classifier $\bm{w}$ (Arjovsky et al., 2019). Invariant Risk Minimization (IRM) proposes to learn $\Phi(\bm{x})$ by merely using invariant features, which yields the same conditional distribution $\mathbb{P}^{t}(\bm{y}|\Phi(\bm{x}))$ across all domains $\bm{t}$, i.e., $\mathbb{P}(\bm{y}|\Phi(\bm{x}))=\mathbb{P}(\bm{y}|\Phi(\bm{x}),\bm{t})$, which is equivalent to $\bm{y}\perp\bm{t}|\Phi(\bm{x})$ (Arjovsky et al., 2019). Existing approximation methods propose to learn a featurizer by matching $\mathbb{E}^{t}[\bm{y}|\Phi(\bm{x})]$ for different $\bm{t}$ (See Section 2.1 for details).
@@
A case study on REx. However, in the continuous environment setting, there are only very limited samples in each environment. So the finite sample estimations $\hat{\mathbb{E}}^{t}[\bm{y}|\Phi(\bm{x})]$ can deviate significantly from the expectation $\mathbb{E}^{t}[\bm{y}|\Phi(\bm{x})]$. In Section 2.2, We conduct a theoretical analysis of REx (Krueger et al., 2021), a popular variant of IRM. Our analysis shows that when there is a large number of domains and limited sample sizes per domain (i.e., in the continuous domain tasks), REx fails to identify invariant features with a constant probability. This is in contrast to the results in discrete domain tasks, where REx can reliably identify invariant features given sufficient samples in each discrete domain.
@@
The generality of our results. Other popular approximations of IRM also suffer from this limitation, such as IRMv1 (Arjovsky et al., 2019), REx (Krueger et al., 2021), IB-IRM (Ahuja et al., 2021), IRMx (Chen et al., 2022), IGA (Koyama & Yamaguchi, 2020), BIRM (Lin et al., 2022a), IRM Game (Ahuja et al., 2020), Fisher (Rame et al., 2022) and SparseIRM (Zhou et al., 2022). They employ positive losses on each environment $\bm{t}$ to assess whether $\hat{\mathbb{E}}^{t}[\bm{y}|\Phi(\bm{x})]$ matches with the other environments. However, since the estimations $\hat{\mathbb{E}}^{t}[\bm{y}|\Phi(\bm{x})]$ are inaccurate, these methods fail to identify invariant features. We then empirically verify the ineffectiveness of these popular invariance learning methods in handling continuous domain problems. One potential naive solution is to split continuous domains into discrete ones. However, this splitting scheme ignores the underlying relationship among domains and therefore can lead to sub-optimal performance, which is further empirically verified in Section 2.2
@@
Figure 1: Illustration of distributional shifts in discrete and continuous domains (Wang et al., 2020). Existing IRM methods focus on discrete domains, which is inconsistent with many real-world tasks. Our work therefore aims to extend IRM to continuous domains.
@@
Our methods. Recall that our task is to learn a feature extractor $\Phi(\cdot)$ that extracts invariant features from $\bm{x}$, i.e., learn $\Phi$ which elicits $\bm{y}\perp\bm{t}|\Phi(\bm{x})$. Previous analysis shows that it is infeasible to align $\mathbb{E}^{t}[\bm{y}|\Phi(\bm{x})]$ among different domains $\bm{t}$ in continuous domain tasks. Instead, we propose to align $\mathbb{E}^{y}[\bm{t}|\Phi(\bm{x})]$ for each class $\bm{y}$. We can start by training two domain regressors, i.e., $h(\Phi(\bm{x}))$ and $g(\Phi(\bm{x}),\bm{y})$, to regress over the continuous domain index $\bm{t}$ using L1 or L2 loss. Notably, the L1/L2 distance between the predicted and ground-truth domains captures the underlying relationship between continuous domain indices and can lead to similar domain index prediction losses between domains on both $h$ and $g$ if $\Phi(\bm{x})$ only extracts invariant features. Specifically, the loss achieved by $h$, i.e., $\mathbb{E}\|\bm{t}-h(\Phi(\bm{x}))\|_{2}^{2}$, and the loss achieved by $g$, i.e., $\mathbb{E}\|\bm{t}-g(\Phi(\bm{x}),\bm{y})\|_{2}^{2}$ would be the same if $\Phi(\bm{x})$ only extracts invariant features. The whole procedure is then formulated into a mini-max framework as shown in Section 3.
@@
In Section 3.1, we emphasize the theoretical advantages of CIL over existing IRM approximation methods in continuous domain tasks. The limited number of samples in each domain makes it challenging to obtain an accurate estimation for $\mathbb{E}^{t}[\bm{y}|\Phi(\bm{x})]$, leading to the ineffectiveness of existing methods. However, CIL does not suffer from this problem because it aims to align $\mathbb{E}^{y}[\bm{t}|\Phi(\bm{x})]$, which can be accurately estimated due to the ample number of samples available in each class (also see Appendix D for the discussion of the relationship between $\mathbb{E}^{t}[\bm{y}|\Phi(\bm{x})]$ and $\mathbb{E}^{y}[\bm{t}|\Phi(\bm{x})]$). In Section 4, we carry out experiments on both synthetic and real-world datasets (including data collected from production systems in Alipay and vision datasets from Wilds-time (Yao et al., 2022)). CIL consistently outperforms existing invariance learning baselines in all the tasks. We summarize our contributions as follows:
@@
  • We identify the problem of learning invariant features under continuous domains and theoretically demonstrate that existing methods that treat domains as categorical indexes can fail with constant probability, which is further verified by empirical study.
@@
  • We propose Continuous Invariance Learning (CIL) as the first general training framework to address this problem. We also show the theoretical advantages of CIL over existing IRM approximation methods on continuous domain tasks.
@@ 2.1 Preliminaries
Notations. Consider the input and output $(\bm{x},\bm{y})$ in the space $\mathcal{X}\times\mathcal{Y}$. Our task is to learn a function $f_{\theta}\in\mathcal{F}:\mathcal{X}\xrightarrow[]{}\mathcal{Y}$, parameterized by $\theta$. We further denote the domain index as $\bm{t}\in\mathcal{T}$. We have $(\bm{x},\bm{y},\bm{t})\sim\mathbb{P}(\bm{x},\bm{y}|\bm{t})\times\mathbb{P}(\bm{t})$. Denote the training dataset as $\mathcal{D}^{\text{tr}}:=\{(\mathbf{x}_{i}^{\text{tr}},\mathbf{y}_{i}^{\text{tr}},\mathbf{t}_{i}^{\text{tr}})\}_{i=1}^{n}$. The goal of domain generalization is to learn a function $f_{\theta}$ on $\mathcal{D}^{\text{tr}}$ that can perform well in unseen testing dataset $\mathcal{D}^{\text{te}}$. (This is different from continuously indexed domain adaptation (Wang et al., 2020), as our setting does not have access to target-domain data.) Since the testing domain $\bm{t}^{\text{te}}$ differs from $\bm{t}^{\text{tr}}$, $\mathbb{P}(\bm{x},\bm{y}|\bm{t}^{\text{te}})$ is also different from $\mathbb{P}(\bm{x},\bm{y}|\bm{t}^{\text{tr}})$ due to distributional shift. Following Arjovsky et al. (2019), we assume that $\bm{x}$ is generated from invariant feature $\bm{x}_{v}$ and spurious feature $\bm{x}_{s}$ by some unknown function $\xi$, i.e., $\bm{x}=\xi(\bm{x}_{v},\bm{x}_{s})$. By the invariance property (more details shown in Appendix C), we have $\bm{y}\perp\bm{t}|\bm{x}_{v}$ for all $\bm{t}\in\mathcal{T}$. The target of invariance learning is to make $f_{\theta}$ only dependent on $\bm{x}_{v}$.
@@
Invariance Learning. To learn invariant features, Invariant Risk Minimization (IRM) first divides the neural network $\theta$ into two parts, i.e., the feature extractor $\Phi$ and the classifier $\bm{w}$ (Arjovsky et al., 2019). The goal of invariance learning is to extract $\Phi(\bm{x})$ which satisfies $\bm{y}\perp\bm{t}|\Phi(\bm{x})$.
@@
Existing Approximation Methods. Suppose we have collected the data from a set of discrete domains, $\mathcal{T}=\{\bm{t}_{1},\bm{t}_{2},...,\bm{t}_{M}\}$. The loss in domain $\bm{t}$ is $\mathcal{R}^{t}(\bm{w},\Phi):=\mathbb{E}^{t}[\ell(\bm{w}(\Phi(\bm{x})),\bm{y})]$ Since is hard to validate $\bm{y}\perp\bm{t}|\Phi(\bm{x})$ in practice, existing works propose to align $\mathbb{E}^{t}[\bm{y}|\Phi(\bm{x})]$ for all $t$ as an approximation. Specifically, if $\Phi(\bm{x})$ merely extract invariant features, $\mathbb{E}^{t}[\bm{y}|\Phi(\bm{x})]$ would be the same in all $\bm{t}$. Let $\bm{w}^{t}$ denote the optimal classifier for domain $\bm{t}$, i.e., $\bm{w}^{t}:=\operatorname*{arg\,min}_{\bm{w}}\mathcal{R}^{t}(\bm{w},\Phi)$. We have $\bm{w}^{t}(\Phi(\bm{x}))=\mathbb{E}^{t}[\bm{y}|\Phi(\bm{x})]$ hold if the function space of $\bm{w}$ is sufficiently large (Li et al., 2022). So if $\Phi(\bm{x})$ relies on invariant features, $\bm{w}^{t}$ would be the same in all $\bm{t}$. Existing approximation methods try to ensure $\bm{w}^{t}$ that is the same for all environments to align $\mathbb{E}^{t}[\bm{y}|\Phi(\bm{x})]$ (Arjovsky et al., 2019; Lin et al., 2022a; Ahuja et al., 2021). Further variants also include checking whether $\mathcal{R}^{t}(\bm{w},\Phi)$ is the same (Krueger et al., 2021; Ahuja et al., 2020; Zhou et al., 2022; Chen et al., 2022), or the gradient is the same for all $\bm{t}$ (Rame et al., 2022; Koyama & Yamaguchi, 2020). For example, REx (Krueger et al., 2021) penalizes the variance of the loss in domains:
$\mathcal{L}_{\text{REx}}=\sum\nolimits_{t\in\mathcal{T}}\mathcal{R}^{t}(\bm{w},\Phi)+\lambda|\mathcal{T}|\text{Var}(\mathcal{R}^{t}(\bm{w},\Phi)).$ (1)
where $\text{Var}(\mathcal{R}^{t}(\bm{w},\Phi))$ is the variance of the losses among domains.
Remark: The REx loss is conventionally $\mathcal{L}_{\text{REx}}=\frac{1}{|\mathcal{T}|}\sum\nolimits_{t\in\mathcal{T}}\mathcal{R}^{t}(\bm{w},\Phi)+\lambda\text{Var}(\mathcal{R}^{t}(\bm{w},\Phi)).$ We re-scale both terms by $|\mathcal{T}|$ for ease of presentation, which does not change the results.